OpenStudy (anonymous):

The sum of the squares of 3 consecutive positive integers is 116.

OpenStudy (kirbykirby):

a+(a+1)+(a+2)=116

OpenStudy (kirbykirby):

oops

OpenStudy (kirbykirby):

a^2+(a+1)^2+(a+2)^2=116

OpenStudy (kirbykirby):

hm it doesn't seem to work though. a is not an integer =\

OpenStudy (anonymous):

These are the options that it gave me 3n2+ 5 = 116 3n2+ 3n + 3 = 116 3n2+ 6n + 5 = 116

OpenStudy (kirbykirby):

3n2+ 6n + 5 = 116

OpenStudy (kirbykirby):

this is weird though since if you solve for x, x won't be an integer...

OpenStudy (kirbykirby):

(well "n" lol sorry i changed variables)

OpenStudy (anonymous):

how did you get that 3n2+6n+5=116? that's what i don't know how to get by myself

OpenStudy (kirbykirby):

\(n^2+(n+1)^2+(n+2)^2\) \(=n^2+(n^2+2n+1)+(n^2+4n+4)\) \(=3n^2+6n+5\)

OpenStudy (kirbykirby):

expand everything and combine like-terms

OpenStudy (anonymous):

oh, haha I didn't see that before lol thanks ^-^

OpenStudy (kirbykirby):

:)